HUBBLE SPACE TELESCOPE DAILY REPORT #1534 PERIOD COVERED: 0000Z (UTC) 12/11/95 - 0000Z (UTC) 12/12/95 Daily Status Report as of 346/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) = The WF/PC-2 was used to make snapshot images of radio galaxy = 3C180.0 as part of a continuum survey and to study the galaxy's = morphology. The observations were executed as scheduled, and no anomalies = were noted. 1.2 Completed FOS/BL 6055 (Spectrometry of the Central W-R Star of the = LMC-PN-N66) The Faint Object Spectrograph (Blue Side) was used to make = observations of the central star in the planetary nebula LMC-N66-CS which = is located in the Large Magellanic Cloud. This star has recently = developed Wolf-Rayet type features and the continuum has brightened = significantly in the ultraviolet and visible. The observations were = executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 6033 (Parallel High Resolution Imaging of Defuse = Objects in the Magellanic Clouds) The WF/PC-2 was used to make parallel images of a variety of defuse = objects located in the Magellanic Clouds. The observations were primarily = in the Balmer lines and stronger forbidden lines, with supplemental = continuum images. The observations were completed as scheduled, and no = problems were reported. 1.4 Completed FGS 6063 (Continuation of Mass Ratio Determination for = Doubles) The Fine Guidance Sensors were used to make POS and TRANS mode = astrometric observations of double star L722-22-00. This binary star = system contains a low mass star and a brown dwarf. The observations were = completed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported. 1.6 Completed FOC/96 5822 (Wind Accretion in the Mira AB System) The Faint Object Camera (f/96) was used to make multi-wave images = and objective prism spectra of the detached binary variable star HD14386 = (aka, the Mira AB System) to study its wind accretion. This system = consists of a compact object and a red giant star. The compact object is = accreting mass from the giant's stellar wind. The observations were = completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration = images. Both photometric and narrow-band filters were used. The = observations were executed as planned, and no problems were reported. 1.8 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.9 Completed FOS/BL 5851 (Post Asymptotic Giant Branch Evolution in = the Magellanic Clouds) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet observations of the planetary nebula LMC-SMP78 which is = located in the Large Magellanic Cloud. Extensive ground-based and HST = observations and imaging work has already been done on this object. The = observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky survey = observations at latitudes larger than 10 degrees taken in parallel with = the prime science observations. The observations were executed as = scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 There was a three minute, 29 second loss of lock starting at = 346/042924Z. The take data flag went down as a SIP was being commanded. Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: HST operations have resumed in the MOR after the completion of the = reconfiguration activities. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH